6 August 2017
Saturday 5th and Sunday 6th of August 2017

The yellow tent camp and view into an Arctic Oven Tent.

EGRIP tent camp

As a new thing this year we have a tent camp for accommodation. The yellow tents are unheated but the volume is smaller so they heat well in the sun and during the night while you are a sleep. The EGRIP hotel thus offers accommodation ranging from the warm and noisy dome, the heated red weatherports and ’Tomatoes’ and the colder yellow tents.

What we did today:

  1. Deep drilling core length at 19.00 Saturday: 8.71 m.
  2. Drillers depth at 19.00 Saturday: 831.46 m.
  3. Borehole logging and calibration and filter runs Sunday.
  4. Shallow drilling S3 ice core Saturday: length 14.03 m, depth 107.89 m.
  5. Water vapour, methane, aerosols, snow and boundary layer measurements.
  6. Physical properties measurements.
  7. Water isotope sampling for sublimation studies.
  8. Grooming of skiway.
  9. Main and Backup generator and Pistenbully maintainance.
  10. White garage tent emptied.
  11. Floor grids up to surface from trenches.
  12. Japanese pit study.
  13. Cleaning, tidying and documenting in the carpenters garage.
  14. Japanese Saturday night dinner - outstanding.

Ad 1-3 Deep drilling this week was relatively stable. We introduced alternating core barrels from one run to the next to allow the core barrel to relax in the logging cabin for at least one hour before extracting the core. Brittle ice core quality is still generally fair/good. Both barrels are fitted with step cutters set with a pitch of about 1.6 mm. Drilling currents are 8-9A for the first 1m or so, and up to 10-11A to the end of our established maximum core length of 1.8 m. We also began alternating hallow shafts between about every two runs to allow more thorough cleaning in the heated box. We replaced the new triple bladed 3-D printed boosters with the classic single bladed boosters on both shafts. We did this after we observed systematic hard packed chips at the booster that would at times cause us to abort runs early due to high current (13A or higher) or runs with poor chips recovery. So far the classic boosters have performed well, and chips recovery has been good. Two filter runs throughout the week produced very little chips from the borehole (less than 3 kg each). An error in orienting the outer core barrel with respect to the Azimuth when remounting drill after the Egg experiment caused a negative trend in inclination, but has since been discovered and corrected and the inclination is again on the way down.
The shallow drill and winch have been performing without trouble. We are using new cutters set to 5 mm pitch, and core quality is still very good at 110m.

Weather: Sunny with periods of overcast with snow showers, wind 1-6 m/s  from N and W, temperatures -20 °C to -8 °C.

FL, Dorthe Dahl-Jensen

← Previous entry   Next entry →